OpenStudy (anonymous):

Fans & Medals!! Simplify: -7(8 – d) Simplify: -7(8 – d) A.-56 + 7d B.-56 + d C.-49d D.49d

OpenStudy (anonymous):

 distribute

OpenStudy (anonymous):

meaning \[-7 \times 8 + 7 \times d\]

OpenStudy (anonymous):

The answer is A because when you distribute -7 X 8 = -56 and -7 X -d = 7d because 2 negatives equal a positive.